I have dice in my hand and want to know what's the probability is of rolling an 8?
please explain...

Answer by checkley77(12844)

2+6=8
3+5=8
4+4=8
5+3=8
6+2=8
5 watys to get a total roll of 8 out of:
6*6=36 total outcomes.
5/36=.13889 or 13.889% in rollling an eight.